OpenStudy (anonymous):

can someone please help me solve -5r+6 greater than or equal to -5(r+2)

hero (hero):

\(-5r + 6 \ge -5(r + 2)\)

OpenStudy (anonymous):

yes thats correct

hero (hero):

Add 5(r + 2) to both sides Add 5r to both sides Subtract 6 from both sides: \(5(r + 2) \ge 5r - 6\)

hero (hero):

Distribute 5 to r and 2: \(5r + 10 \ge 5r - 6\) Subtract 5r from both sides: Subtract 10 from both sides: \(5r - 5r \ge -6 - 10\) \(0 \ge -16 \) False

OpenStudy (anonymous):

ok thanks but it was actually less than or equal to so im sorry about that i got a little confused @Hero

hero (hero):

Either way there's still no solution.